# Collective excitations in chiral spin liquid: chiral roton and long-wavelength nematic mode

### v1

@@ Abstract
Chiral spin liquid (CSL) is a magnetic analogue of the fractional quantum Hall (FQH) liquid. Collective excitations play a vital role in shaping our understanding of these exotic quantum phases of matter and their quantum phase transitions. While the magneto-roton and long-wavelength chiral graviton modes in the FQH liquids have been extensively explored, the collective excitations of CSLs remain elusive. Here we explore the collective excitations in the SU(2) symmetric CSL phase of the spin-1/2 square-lattice $J_{1}-J_{2}-J_{\chi}$ model, where an intriguing quantum phase diagram was recently revealed. Combining exact diagonalization and time-dependent variational principle calculations, we observe two spin-singlet collective modes: a chiral p-wave (low-energy) roton mode at finite momentum and a d-wave (higher-energy) nematic mode at zero momentum, both of which are prominent across the CSL phase. Such exotic modes exhibit fingerprints distinct from those of FQH liquids, and to the best of our knowledge, are reported for the first time. By tuning $J_{2}$, we find the nematic mode to be pronouncedly soft, together with the spin-triplet two-spinon bound states, potentially promoting strong nematic and spin stripe instabilities. Our work paves the way for further understanding CSL from the dynamical perspective and provides new spectroscopic signatures for future experiments of CSL candidates.
@@
In FQH systems, the prototypical example of a neutral collective excitation is the magneto-roton mode, normally with energy minima at finite wave vectors Girvin et al. (1985, 1986); Yang (2013). In the long-wavelength limit, a distinct collective mode emerges at higher energy: a chiral and quadrupolar orbital spin-2 excitation, also known as the chiral graviton mode Haldane (2011a); Yang et al. (2012); Golkar et al. (2016). These modes have been numerically identified Liu et al. (2018); Liou et al. (2019); Kumar and Haldane (2022); Liu et al. (2024) and experimentally detected in two-dimensional electron gas Kukushkin et al. (2009); Liang et al. (2024). And their nature in fractional Chern insulators (FCIs) – the zero-field lattice-based cousins of FQH states – has also spurred intensive theoretical investigations Repellin et al. (2014); Lu et al. (2024a); Kousa et al. (2025); Long et al. (2025); Wang et al. (2025); Shen et al. (2025); Long et al. (2026). CSLs share profound similarities with FQH liquids in their ground-state topology. Yet, it remains unexplored whether — and in what form — their collective excitations exhibit analogous phenomenology, particularly in SU(2)-invariant quantum magnets.
Figure 1: Lattice model and phase diagram. (a) The $J_{1}-J_{2}-J_{\chi}$ model on square lattice. (b) Schematic plot of the two exotic spin-singlet collective modes deep inside the CSL phase. For positive $J_{\chi}$ the chiral roton mode lies in the $p+\mathrm{i}p$ channel. (c) The quantum phase diagram with fixed $J_{1}=1,\ J_{2}=0$ and tuning $J_{\chi}$. The GS in the shaded region is still under debate and is not the focus of this work.
@@
Here, we address these open questions by systematically investigating the collective excitations of a SU(2) symmetric CSL. Intriguingly, we discover a chiral p-wave roton mode and a d-wave nematic mode in the spin-singlet sector [Fig. 1(b)], both of which are distinct from the known knowledge of the FQH liquid. These exotic collective modes are directly revealed, combining exact diagonalization (ED), density matrix renormalization group (DMRG), and time-dependent variational principle (TDVP) calculations on the spin-1/2 $J_{1}-J_{2}-J_{\chi}$ model on square lattice. The p-wave roton is the lowest energy mode, found at $\mathbf{q}=(\pi,\pi)$, with chirality locked to the GS chirality, and it remains prominent even near the phase boundaries. The higher-energy d-wave mode at $\mathbf{q}=0$ can be pronouncedly softened at finite $J_{2}$, indicating the nematic instability. In addition, the spin-triplet collective excitations are also studied, with some finite-momentum two-spinon bound states observed. These findings offer insights into the dynamical landscape of CSL, thereby prompting a fundamental question concerning the nature of CSL: how different is CSL from the established FQH/FCI paradigm despite their ground states are topologically equivalent?
@@
$\mathrm{DSF}^{\hat{O}}(\mathbf{q},\omega)=\sum_{\mathbf{r},t}\frac{e^{\mathrm{i}\omega t-\eta t}e^{-\mathrm{i}\mathbf{q}\mathbf{r}}}{\sqrt{N_{t}}}(\langle\hat{O}^{\dagger}_{\mathbf{r}}(t)\hat{O}_{\mathbf{0}}(0)\rangle-\langle\hat{O}^{\dagger}_{\mathbf{r}}\rangle\langle\hat{O}_{\mathbf{0}}\rangle),$ (2)
@@
First, we begin with the triplet sector, as is customary in the study of spin liquids, and present the spin DSFs by targeting $\hat{O}=S^{z}$ operator in Fig. 2(a1-a3), where the full dispersion along $q_{x}$ is resolved at several relevant $q_{y}$. Rather than a purely featureless continuum, we find some finite-momentum bound states with relatively concentrated spectral weight. One notable feature is that these modes appear in relatively high-frequency regime.
@@
Nature of the chiral roton.— To provide further insight on the observed roton mode, we show the energy spectrum of the CSL on torus in Fig.3. We find that the lowest excitation belongs to the spin singlet and is located at $(\pi,\pi)$. Its quantum number matches that of the roton mode we found in Fig. 2(b-c). The precise nature of this low-energy excitation can be obtained by inspecting its content using the SMA. As shown in Tab.1, on the symmetric torus, we find very high overlaps between the target excited state and the variational states made of $\hat{B}_{x}$ and $\hat{B}_{y}$ ove . More interestingly, we find the relation $\hat{B_{x}}|\tilde{\psi}_{0}(0,0)\rangle=\mathrm{i}\hat{B_{y}}|\tilde{\psi}_{0}(0,0)\rangle$ exactly holds. This motivates us to design the chiral p-wave operators $\hat{P}^{\pm}=\hat{B}_{x}\pm\mathrm{i}\hat{B}_{y}$. The overlap of the new chiral states is either close to $1$ (larger than the $\hat{B}_{\mu}$ operators) or 0, and we have checked that the chirality is locked to the GS chirality, which could be simply flipped by the sign of $J_{\chi}$. We have also checked that the chirality of this excited state is well retained even for the non-symmetric torus such as the $4\times 6$ one [Tab. 1], and the aforementioned properties are consistently found across the CSL phase sup .
@@
Figure 4: The chiral roton. DSFs of the chiral p-wave operators $\hat{P}^{\mp}$ at fixed $\omega=0.28$ (energy scale of the roton mode) for $J_{2}=0$ and $J_{\chi}=2$ are shown, respectively. The calculations are performed by the TDVP on $L_{y}\times L_{x}=4\times 32$ cylinders. The black dashed lines denote $q_{y}=\frac{\pi}{2},\pi,\frac{3\pi}{2}$, respectively.
@@
Furthermore, we test the SMA picture by computing the static structure factors (SSFs). For the $\hat{P}^{+}$ channel, there is a peak at ($\pi,\pi$), which is stronger than those of the $\hat{B}_{\mu}$ operators even after renormalization sup , implying the larger frequency-integrated dynamical spectral weight in this chiral channel. Interestingly, for the $\hat{P}^{-}$ channel, the SSF at ($\pi,\pi$) is $\sim 0$. These results further support that the low-energy state at ($\pi,\pi$) we observe is consistent with a chiral p-wave mode, which we call the chiral roton mode. To further confirm this chiral p-wave roton mode, we show the DSFs of the $\hat{P}^{\pm}$ channels at fixed $\omega=0.28$ (the energy of the roton mode) from TDVP simulations in Fig. 4. The full dispersion is included in the SI sup . We observe a sharp peak at $(\pi,\pi)$ in the $\hat{P}^{+}$ channel, while there is no response in the $\hat{P}^{-}$ one. These are consistent with the earlier ED results in Fig. 3 and Tab. 1, and together support the chiral p-wave nature of the observed singlet roton mode. This unexpected discovery distinguishes this CSL from the Laughlin FQH states whose magnetoroton modes at the finite-wave-vector minima do not carry any independent chirality quantum number, since it is created in the scalar density channel, despite the GS chirality Girvin et al. (1986); Kumar and Haldane (2022).
@@
Interestingly, we find that the $\hat{D}_{x^{2}-y^{2}}$ operator exhibits a one-to-one (diagonal) selection rule between the two degenerate GSs and the two lowest excited states (circled in blue in Fig.5(a)) with very high overlaps, supporting that these excitations are well captured by a single-mode ansatz. In addition, according to the overlaps of the $\hat{D}^{\pm}$ channels, no obvious chirality is observed. We further show the DSFs of different d-wave operators with fixed $q_{y}=0$ on the $4\times 32$ cylinder in Fig.5 (b-d) sup . It is consistent with the ED results that the nematic mode (at higher energy at parameters deep in the CSL phase) becomes pronouncedly soft, and the DSF of the nematic mode is significantly sharp with highly concentrated spectral weight, exhibiting strong nematic instability. Although the different chiral d-wave operators show different absorption strengths, they are of the same order of magnitude, and the peak positions remain essentially unchanged.
@@
Discussions and implications.— In this work, we have presented a systematic study of the collective excitations of CSL. Although the topological order of many-body ground states is the same as the 1/2 Laughlin FQH/FCI states, we find their low-lying collective excitations are quite different. In the singlet sector, we observe a prominent low-energy chiral p-wave roton mode at finite momentum, which is prominent throughout the phase parameter space. By contrast, in the FQH/FCI systems, the roton minimum of the magnetoroton mode exhibits no chirality or orbital spin Girvin et al. (1986). This is an unexpected result and is reported for the first time, to the best of our knowledge. Furthermore, we have also observed a $\mathbf{q}=0$ d-wave nematic mode, which is similar to the chiral graviton mode in FQH/FCI systems in the aspect of the orbital spin-2 and the fact that it is at higher energy when far from the phase boundaries Liou et al. (2019). However, we observe no obvious chirality of the nematic mode in the CSL. Although the chiral roton mode and the nematic mode exhibit different characteristics, the dispersions suggest that they may be connected. Whether they belong to the same excitation branch is interesting for further studies.
@@
Our findings underscore the significance of investigating the collective excitations in the CSLs, since these systems are often assumed to exhibit only a continuum. The collective modes could be intrinsic to such topological liquids and might play a role comparable to that of the magnetoroton and graviton modes in shaping our understanding of FQH liquids Girvin et al. (1986); Haldane (2011b). This could also motivate further investigations of collective modes in other CSLs (such as non-abelian ones) and different QSLs (with different or without topological orders). Since the model in this work is based on the square lattice with $C_{4}$ symmetry, it would be interesting for future work to explore how the lattice geometries and symmetries could affect the collective modes discovered in this work. Note that, this could be very different from the Laughlin FCIs with different discrete rotation symmetries, where the magnetoroton minima are consistently found without orbital spin or chirality, and the spin-2 gravitons are consistently found to be chiral Repellin et al. (2014); Long et al. (2025, 2026); Wang et al. (2025); Kousa et al. (2025).
@@
Experimentally, detecting a CSL is often more demanding than detecting a fermionic FQH/FCI phase. In FQH/FCI systems, topological order is accompanied by quantized charge transport responses Stormer et al. (1999), while in CSLs the corresponding signatures are carried by neutral spin/energy currents and are typically accessed only indirectly Haldane and Arovas (1995); Zhang et al. (2024b). Therefore, additional experimental probes are highly desirable, and one promising route is to detect the dynamical properties and excitations. Inelastic neutron scattering is the primary momentum-resolved probe of spin excitations, and it predominantly accesses the $S=1$ channel in an SU(2)-symmetric system with a singlet ground state Broholm et al. (2020). However, a broad continuum and even the bound-state-like features in the triplet channel are insufficient for identification. In this work, the singlet collective modes offer new spectroscopic signatures for future experiments on candidate CSL materials. For example, the $\mathbf{q}=0$ nematic mode could be identified from the (circularly) polarized Raman scattering measurements Devereaux and Hackl (2007); Liang et al. (2024); Koller et al. (2025), where symmetry selection rules allow one to isolate the d-wave channel ($B_{1g}$ channel for this model) Sandvik et al. (1998), and it is also interesting to experimentally explore whether the nematic mode is chiral. A possible route to access the finite-$\mathbf{q}$ chiral p-wave roton mode is momentum- and polarization-resolved (indirect) resonant inelastic X-ray scattering measurements Ament et al. (2011); Sikora et al. (2010), where circular-polarization dependence might provide a direct way to distinguish $\hat{P}^{\pm}$ channels through chiral selection rules.
@@
We close by highlighting the soft collective modes in the CSL are related to interesting quantum phase transitions. For example, we study the soft modes at $J_{2}=0.6$ and $J_{\chi}=0.5$, but the GS phase diagram at fixed $J_{\chi}=0.5$ larger $J_{2}$ is still under debate. Ref. Zhang et al. (2024a) suggests a direct CSL-stripe transition at $J_{2}\sim 0.75$ based on DMRG results and the spin stripe order is ($0,\pi$) or ($\pi,0$), However, Ref. Yang et al. (2024) suggests the possibility of an intermediate nematic spin liquid phase between the CSL and stripe phases, based on the 32-site ED results, but the identification of the nematic phase still suffers from the finite-size effect. In fact, the soft collective modes reported in our work are consistent with both scenarios. For the first one, if the CSL-stripe transition is continuous, one might observe the $\mathbf{q}=0$ nematic mode and ($0,\pi$)/($\pi,0$) triplet mode both become gapless at the critical point, and the stripe phase itself has the co-existing secondary nematic order. For the second scenario, the nematic spin liquid phase shall be a nematic CSL, and the two-step transition could be triggered by the sequential gap closing of the nematic and triplet modes in the CSL. This two-step transition also leads to a unique vestigial scenario Nie et al. (2017); Fernandes et al. (2019); Lu et al. (2024b), in which the vestigial nematic phase inherited from stripe order is topologically ordered. The mechanism for the nematic CSL discussed here echoes that of FQH nematics Regnault et al. (2017); Yang (2020); Pu et al. (2024) and smectic FCIs Lu et al. (2024c, 2025), where the latter phases are triggered by the soft chiral graviton at $\mathbf{q}=0$ and magneto-roton at finite momentum, respectively, and the topological orders remain intact through the symmetry-breaking process. A more extensive exploration of the parameter space is required, and either interesting scenario may be realized in a broader regime, which we leave for future work.

### v2

@@ Abstract
Chiral spin liquid (CSL) is a magnetic analogue of the fractional quantum Hall (FQH) liquid. Collective excitations play a vital role in shaping our understanding of these exotic quantum phases of matter and their quantum phase transitions. While the magneto-roton and long-wavelength chiral graviton modes in the FQH liquids have been extensively explored, the collective excitations of CSLs remain elusive. Here we explore the collective excitations in the SU(2) symmetric CSL phase of the spin-1/2 square-lattice $J_{1}-J_{2}-J_{\chi}$ model, where an intriguing quantum phase diagram was recently revealed. Combining exact diagonalization and time-dependent variational principle calculations, we observe two spin-singlet collective modes: a chiral p-wave (low-energy) roton mode at finite momentum and a elliptically polarized d-wave (higher-energy) nematic mode at zero momentum, both of which are prominent across the CSL phase. Such exotic modes exhibit fingerprints distinct from those of FQH liquids, and to the best of our knowledge, are reported for the first time. By tuning $J_{2}$, we find the nematic mode to be pronouncedly soft, together with the spin-triplet two-spinon bound states, potentially promoting strong nematic and spin stripe instabilities. Our work paves the way for further understanding CSL from the dynamical perspective and provides new spectroscopic signatures for future experiments of CSL candidates.
@@
In FQH systems, the prototypical example of a neutral collective excitation is the magneto-roton mode, normally with energy minima at finite wave vectors Girvin et al. (1985, 1986); Yang (2013). In the long-wavelength limit, a distinct collective mode emerges at higher energy: a chiral and quadrupolar orbital spin-2 excitation, also known as the chiral graviton mode Haldane (2011a); Yang et al. (2012); Golkar et al. (2016). These modes have been numerically identified Liu et al. (2018); Liou et al. (2019); Kumar and Haldane (2022); Liu et al. (2024) and experimentally detected in two-dimensional electron gas Kukushkin et al. (2009); Liang et al. (2024). And their nature in fractional Chern insulators (FCIs) – the lattice-based cousins of FQH states – has also spurred intensive theoretical investigations Repellin et al. (2014); Lu et al. (2024a); Kousa et al. (2025); Long et al. (2026a); Wang et al. (2025); Shen et al. (2025); Xavier et al. (2025); Long et al. (2026b). CSLs share profound similarities with FQH liquids in their ground-state topology. Yet, it remains unexplored whether — and in what form — their collective excitations exhibit analogous phenomenology, particularly in SU(2)-invariant quantum magnets.
Figure 1: Lattice model and phase diagram. (a) The $J_{1}-J_{2}-J_{\chi}$ model on square lattice. (b) Schematic plot of the two exotic spin-singlet collective modes deep inside the CSL phase. For positive $J_{\chi}$ the chiral roton mode lies in the $p+\mathrm{i}p$ channel, and the elliptically polarized nematic mode exhibits stronger $d+\mathrm{i}d$ component. (c) The quantum phase diagram with fixed $J_{1}=1,\ J_{2}=0$ and tuning $J_{\chi}$. The GS in the shaded region is still under debate and is not the focus of this work.
@@
Here, we address these open questions by systematically investigating the collective excitations of a SU(2) symmetric CSL. Intriguingly, we discover a chiral p-wave roton mode and a elliptically polarized d-wave nematic mode in the spin-singlet sector [Fig. 1(b)], with distinc features from the known knowledge of the FQH liquid. These exotic collective modes are directly revealed, combining exact diagonalization (ED), density matrix renormalization group (DMRG), and time-dependent variational principle (TDVP) calculations on the spin-1/2 $J_{1}-J_{2}-J_{\chi}$ model on square lattice. The p-wave roton is the lowest energy mode, found at $\mathbf{q}=(\pi,\pi)$, with chirality locked to the GS chirality, and it remains prominent even near the phase boundaries. The higher-energy d-wave mode at $\mathbf{q}=0$ is elliptically polarized (the imbalance between the $d\pm\mathrm{i}d$ channels tracks the GS chirality), and can be substantially softened at finite $J_{2}$, indicating the nematic instability. In addition, the spin-triplet collective excitations are also studied, with some finite-momentum two-spinon bound states observed. These findings offer insights into the dynamical landscape of CSL, thereby prompting a fundamental question concerning the nature of CSL: how different is CSL from the established FQH/FCI paradigm despite their ground states are topologically equivalent?
@@
$\mathrm{DSF}^{\hat{O}}(\mathbf{q},\omega)=\sum_{\mathbf{r},t}\frac{e^{\mathrm{i}\omega t-\eta t}e^{-\mathrm{i}\mathbf{q}\mathbf{r}}}{\sqrt{N_{t}}}(\langle\hat{O}^{\dagger}_{\mathbf{r}}(t)\hat{O}_{\mathbf{0}}(0)\rangle-\langle\hat{O}^{\dagger}_{\mathbf{r}}\rangle\langle\hat{O}_{\mathbf{0}}\rangle)$ (2)
@@
First, we begin with the triplet sector, as is customary in the study of spin liquids Zhu et al. (2019); Pichler et al. (2024); Willsher and Knolle (2025), and present the spin DSFs by targeting $\hat{O}=S^{z}$ operator in Fig. 2(a1-a3), where the full dispersion along $q_{x}$ is resolved at several relevant $q_{y}$. Rather than a purely featureless continuum, we find some finite-momentum bound states with relatively concentrated spectral weight. One notable feature is that these modes appear in relatively high-frequency regime.
@@
Nature of the chiral roton.— To provide further insight on the observed roton mode, we show the energy spectrum of the CSL on torus in Fig.3. We find that the lowest excitation belongs to the spin singlet and is located at $(\pi,\pi)$. Its quantum number matches that of the roton mode we found in Fig. 2(b-c). The precise nature of this low-energy excitation can be obtained by inspecting its content using the SMA. As shown in Tab.1, on the symmetric torus, we find very high overlaps between the target excited state and the variational states made of $\hat{B}_{x}$ and $\hat{B}_{y}$ ove . More interestingly, we find the relation $\hat{B_{x}}|\tilde{\psi}_{0}(0,0)\rangle=\mathrm{i}\hat{B_{y}}|\tilde{\psi}_{0}(0,0)\rangle$ exactly holds. This motivates us to design the chiral p-wave operators $\hat{P}^{\pm}=\hat{B}_{x}\pm\mathrm{i}\hat{B}_{y}$. The overlap of the new chiral states is either close to $1$ (larger than the $\hat{B}_{\mu}$ operators) or exactly 0, and we have checked that the chirality is locked to the GS chirality, which could be simply flipped by the sign of $J_{\chi}$. We have also checked that the chirality of this excited state is well retained even for the non-symmetric torus such as the $4\times 6$ one [Tab. 1], and the aforementioned properties are consistently found across the CSL phase sup .
@@
Figure 4: The chiral roton. DSFs of the chiral p-wave operators $\hat{P}^{\mp}$ at fixed $\omega=1.76$ (energy scale of the roton mode) for $J_{2}=0$ and $J_{\chi}=2$ are shown, respectively. The calculations are performed by the TDVP on $L_{y}\times L_{x}=4\times 32$ cylinders. The black dashed lines denote $q_{y}=\frac{\pi}{2},\pi,\frac{3\pi}{2}$, respectively.
@@
Furthermore, we test the SMA picture by computing the static structure factors (SSFs). For the $\hat{P}^{+}$ channel, there is a peak at ($\pi,\pi$), which is stronger than those of the $\hat{B}_{\mu}$ operators even after renormalization sup , implying the larger frequency-integrated dynamical spectral weight in this chiral channel. Interestingly, for the $\hat{P}^{-}$ channel, the SSF at ($\pi,\pi$) is $\sim 0$. These results further support that the low-energy state at ($\pi,\pi$) we observe is consistent with a chiral p-wave mode, which we call the chiral roton mode. To further confirm this chiral p-wave roton mode, we show the DSFs of the $\hat{P}^{\pm}$ channels at fixed $\omega=1.76$ (the energy of the roton mode) from TDVP simulations in Fig. 4. The full dispersion is included in the SI sup . We observe a sharp peak at $(\pi,\pi)$ in the $\hat{P}^{+}$ channel, while there is no response in the $\hat{P}^{-}$ one. These are consistent with the earlier ED results in Fig. 3 and Tab. 1, and together support the exactly chiral p-wave nature of the observed singlet roton mode. This unexpected discovery distinguishes this CSL from the Laughlin FQH states whose magnetoroton modes at the finite-wave-vector minima do not carry any independent chirality quantum number, despite the GS chirality Girvin et al. (1986).
@@
Interestingly, we find that the $\hat{D}_{x^{2}-y^{2}}$ operator exhibits a one-to-one (diagonal) selection rule between the two degenerate GSs and the two lowest excited states (circled in blue in Fig.5(a)) with very high overlaps, supporting that these excitations are well captured by a single-mode ansatz. In addition, according to the overlaps of the $\hat{D}^{\pm}$ channels, no obvious chirality is observed. We further show the DSFs of different d-wave operators with fixed $q_{y}=0$ on the $4\times 32$ cylinder in Fig.5 (b-d) sup . It is consistent with the ED results that the nematic mode (at higher energy at parameters deep in the CSL phase) becomes markedly soft, and the DSF of the nematic mode is significantly sharp with highly concentrated spectral weight, exhibiting strong nematic instability. Furthermore, instead of a pure chirality eigenmode, we find that this d-wave nematic mode is elliptically polarized, as the DSF in the $d+\mathrm{i}d$ channel is approximately twice that in the $d-\mathrm{i}d$ channel. By tuning the relative weight of $\hat{D}_{x^{2}-y^{2}}$ and $\hat{D}_{xy}$ components in the $\hat{D}^{\pm}$ operators, the absorption strength in the $d\pm\mathrm{i}d$ channels could be even more comparable.
@@
Discussions and implications.— In this work, we have presented a systematic study of the collective excitations of CSL. Although the topological order of many-body ground states is the same as the 1/2 Laughlin FQH/FCI states, we find their low-lying collective excitations are quite different. In the singlet sector, we observe a prominent low-energy chiral $p+\mathrm{i}p$ roton mode at finite momentum, which is prominent throughout the phase parameter space. By contrast, in the FQH/FCI systems, the roton minimum of the magnetoroton mode exhibits no chirality or orbital spin Girvin et al. (1986). This is an unexpected result and is reported for the first time, to the best of our knowledge. Furthermore, we have also observed a $\mathbf{q}=0$ d-wave nematic mode, which is similar to the chiral graviton mode in FQH systems in the aspect of the orbital spin-2 and the fact that it is at higher energy when far from the phase boundaries Liou et al. (2019). However, this nematic mode in the CSL is only elliptically polarized: the response is stronger in the $d+\mathrm{i}d$ channel, yet remains comparable to that in the $d-\mathrm{i}d$ channel. Although the chiral roton mode and the elliptically polarized nematic mode exhibit different characteristics, the dispersions suggest that they may be connected. Whether they belong to the same excitation branch is interesting for further studies.
@@
Our findings underscore the significance of investigating the collective excitations in the CSLs, since these systems are often assumed to exhibit only a continuum. The collective modes could be intrinsic to such topological liquids and might play a role comparable to that of the magnetoroton and graviton modes in shaping our understanding of FQH liquids Girvin et al. (1986); Haldane (2011b). This could also motivate further investigations of collective modes in other CSLs (such as non-abelian ones) and different QSLs (with different or without topological orders). Since the model in this work is based on the square lattice with $C_{4}$ symmetry, it would be interesting for future work to explore how the lattice geometries and symmetries could affect the collective modes discovered in this work. Note that, this could be different from the Laughlin FCIs with different discrete rotation symmetries, where the magnetoroton minima are consistently found without orbital spin or chirality, and the spin-2 gravitons are consistently found to be chiral Repellin et al. (2014); Long et al. (2026a, b); Wang et al. (2025); Kousa et al. (2025).
@@
Experimentally, detecting a CSL is often more demanding than detecting a fermionic FQH/FCI phase. In FQH/FCI systems, topological order is accompanied by quantized charge transport responses Stormer et al. (1999), while in CSLs the corresponding signatures are carried by neutral spin/energy currents and are typically accessed only indirectly Haldane and Arovas (1995); Zhang et al. (2024b). Therefore, additional experimental probes are highly desirable, and one promising route is to detect the dynamical properties and excitations. Inelastic neutron scattering is the primary momentum-resolved probe of spin excitations, and it predominantly accesses the $S=1$ channel in an SU(2)-symmetric system with a singlet ground state Broholm et al. (2020). However, a broad continuum and even the bound-state-like features in the triplet channel are insufficient for identification. In this work, the singlet collective modes offer new spectroscopic signatures for future experiments on candidate CSL materials. For example, the $\mathbf{q}=0$ nematic mode could be identified from the (circularly) polarized Raman scattering measurements Sandvik et al. (1998); Devereaux and Hackl (2007); Liang et al. (2024); Koller et al. (2025), where symmetry selection rules allow one to isolate the d-wave channel, and it is also interesting to experimentally explore the (elliptical) polarization of the nematic mode. A possible route to access the finite-$\mathbf{q}$ chiral p-wave roton mode is momentum- and polarization-resolved (indirect) resonant inelastic X-ray scattering measurements Ament et al. (2011); Sikora et al. (2010), where circular-polarization dependence might provide a direct way to distinguish $\hat{P}^{\pm}$ channels through chiral selection rules.
@@
We close by highlighting the soft collective modes in the CSL are related to interesting quantum phase transitions. For example, we study the soft modes at $J_{2}=0.6$ and $J_{\chi}=0.5$, but the GS phase diagram at fixed $J_{\chi}=0.5$ larger $J_{2}$ is still under debate. Ref. Zhang et al. (2024a) suggests a direct CSL-stripe transition at $J_{2}\sim 0.75$ based on DMRG results and the spin stripe order is ($0,\pi$) or ($\pi,0$), However, Ref. Yang et al. (2024) suggests the possibility of an intermediate nematic spin liquid phase between the CSL and stripe phases, based on the 32-site ED results, but the identification of the nematic phase still suffers from the finite-size effect. In fact, the soft collective modes reported in our work are consistent with both scenarios. For the first one, if the CSL-stripe transition is continuous, one might observe the $\mathbf{q}=0$ nematic mode and ($0,\pi$)/($\pi,0$) triplet mode both become gapless at the critical point, and the stripe phase itself has the co-existing secondary nematic order. For the second scenario, the nematic spin liquid phase shall be a nematic CSL, and the two-step transition could be triggered by the sequential gap closing of the nematic and triplet modes in the CSL. This two-step transition also leads to a unique vestigial scenario Nie et al. (2017); Fernandes et al. (2019); Lu et al. (2026), in which the vestigial nematic phase inherited from stripe order is topologically ordered. The mechanism for the nematic CSL discussed here echoes that of FQH nematics Regnault et al. (2017); Yang (2020); Pu et al. (2024) and smectic FCIs Lu et al. (2024b, 2025), where the latter phases are triggered by the soft chiral graviton at $\mathbf{q}=0$ and magneto-roton at finite momentum, respectively, and the topological orders remain intact through the symmetry-breaking process. A more extensive exploration of the parameter space is required, and either interesting scenario may be realized in a broader regime, which we leave for future work.